# CADS: A Systematic Literature Review on the Challenges of Abstractive Dialogue Summarization

### v2

@@ 3.1 Language
Characteristics. The language challenge describes the idiosyncratic nature of spoken language and individual speech patterns. This includes linguistic subtleties such as informal expressions (e.g., ’yeah,’), ungrammatical structures, colloquialisms, personal vocabulary, and linguistic noise (e.g., filler words) (?, ?, ?, ?, ?). It further covers content repetition (i.e., speakers restate or rephrase information for emphasis or clarity) (?, ?, ?), and domain-specific terminology (e.g., medical terms) (?, ?, ?, ?). These elements demand that models accurately interpret and adapt to the linguistic features of dialogues (?). Figure 3 presents a dialogue snippet showing examples of these three sub-challenges. Failing to handle this challenge may result in a loss of coherence and clarity, repetition of content, and factuality issues.
Figure 3: Example dialogue to demonstrate examples of the idiosyncratic nature of spoken language and individual speech patterns. Red displays disfluencies, blue highlights personal speech patterns, orange displays the colloquialism, green represent informal expressions
Approaches. Transformer-based models often struggle with the language characteristics due to a gap between their pre-training data, typically well-edited texts (e.g., news articles, research papers, Wikipedia entries) (?), and the characteristics of spoken dialogue not reflected in these edited texts (?). The transfer between these text styles is hindered by the scarcity of diverse, dialogue-oriented pre-training data, resulting in a lack of exposure to dialogue data during training, consequently reducing models’ performance. Repetition of already provided information is less of an issue for current language models (?).
Current research aims to bridge this gap between the formal language of the pre-training data and dialogues by retraining models with dialogue-focused tasks, considering both single- and multi-task setups. Explored tasks span masking key dialogue elements (e.g., pronouns entities, high-content tokens, and words) (?) and part-of-speech tagging (?). More recent tasks aim to transform informal, first-person dialogue into a structured, third-person narrative through changing speaker names, adjusting grammar, and adding emotional context (?), or simulating a dialogue by modifying pre-training documents into a conversation structure (?). Teaching models to understand the ’who-did-what’ structure through a specific task is further gaining traction (?). Other explored strategies include pre-processing the input for anaphora resolution (?) and adapted pre-training (?, ?, ?).
@@ 3.2 Structure
Characteristics. Aiming at extracting the built-up of a conversation, the structure challenge revolves around segmenting a dialogue transcript and creating an ordered summary based on the extracted structures (?). This challenge involves identifying the topic flow within a conversation (?, ?, ?, ?), tracking the dialogue phases (e.g., problem identification, decision making) (?, ?)), and analyzing the utterance interaction (i.e., relationship and dependencies between utterances) (?, ?, ?). In Figure 4, we show a dialogue snippet exhibiting topic flow, dialogue phases, and utterance interactions. This can result in summaries lacking coherence, completeness, and depth (?, ?, ?, ?, ?).
@@
Approaches. Transformer-based models face two primary hurdles when handling the structural challenge, i.e., adapting to the variability in dialogue structures, such as differing topic sequences and conversation phases depending on the conversation type, and managing long-range dependencies to link utterances. The first difficulty is related to the lack of varied datasets across different dialogue types on which models can train to improve generalization as well as that most pre-trained models used are expecting clearer structural signals due to the more structured data used during pre-training (Section 3.1). The second problem relates to a weakness of transformed-based architectures: their limitation in handling long sequences, which stems from the practical limitation of a fix-sized context window (?), the diluted attention on a single token when many tokens are processed (?), and the bias towards recent information (?).
Recent methods employ graph structures to understand and use dialogue structures more effectively. They combine static and dynamic graphs for a detailed examination of conversation dynamics (?) where the static graphs represent unchanging aspects like speaker relationships, and dynamic graphs track how dialogues evolve. Abstract Meaning Representation (AMR) graphs are employed to capture overarching themes (?), detailed sentence-level connections, and entity interactions, enhancing content comprehension (?). Meanwhile, a hybrid approach combining traditional language models with graph neural networks captures the utterance interactions and addresses the long-range dependency sub-challenge (?). Similarly, ConceptNet (?) is used for thematic linking (?).
To manage long dialogue transcripts beyond context-size limits and the inherent long-distance relations, researchers are moving away from simply truncating inputs to context length to avoid losing vital information. A leading method segments the summarization process into stages (e.g., topics and dialogue phases) for coherent processing (?, ?, ?). The segmentation and multi-pass strategy iteratively summarizes parts of the dialogue and combines partial summaries with a subset of the remaining content in each iteration to build a complete summary (?). The $Summ^{N}$ framework (?), summarizes dialogue segments independently before refining them together, offering a detailed and scalable solution.
Other techniques involve additional training tasks (?, ?, ?, ?), sentence importance measures to group sentences into sub-topics (?, ?), while hierarchical setups to differentiate dialogue parts based on who’s speaking (word level) and what’s being discussed (turn level) are also popular (?).
@@ 3.3 Comprehension
Characteristics. The comprehension challenge involves accurately understanding and contextualizing utterances so models can grasp and leverage these insights for summarization (?). This includes direct and implied content, e.g., unmentioned background knowledge, organizational knowledge, and prior discussions not explicitly referred to (examples in Figure 5). To infer and understand this content, the whole context of a dialogue is required, spanning both the local context, focusing on neighboring words and sentences, and the global context, covering the entire dialogue’s informational flow (?).
Figure 5: Example conversation to demonstrate the comprehension challenge including direct and implied content. Directly stated (orange background) are the explicit updates by Bob and Charlie about their respective tasks. Green marks implied content such as a deadline for a demo and a current focus on performance and finalizing features (unmentioned background knowledge), Redis implies an understanding of its role in performance enhancement (organizational knowledge and reference to prior discussions on data requirements).
Approaches. Transformer-based models struggle with comprehension and contextualization due to the underlying dependency parsing challenges, particularly with unclear sentence boundaries as they appear in spoken language, which can lead to catastrophic forgetting and result in instability in handling long-range dependencies (?). Additionally, these models often interpret the text literally, affecting their ability to grasp nuanced meanings (?, ?). As these models often rely on syntactic relationships between words, they miss the nuanced understanding humans use to interpret implied meanings of sarcasm, irony, euphemisms, and similar pragmatic elements (?). Consequently, this can lead to summaries that miss deeper meanings or implications, resulting in content that may be shallow or misleading (?).
@@ 3.4 Speaker
Characteristics. Participants, their interactions, and how they influence the dynamics of a conversation are part of the speaker challenge. This includes the individual participant roles (e.g., ”project manager”, ”customer service agent”) (?, ?), topic-dependent role changes (?, ?, ?, ?, ?, ?, ?), their interactions (e.g., contributions depending on viewpoints (?, ?, ?), backgrounds (?, ?, ?)), and references to entities and other participants (?, ?, ?, ?). Examples of the speaker characteristics are shown in Figure 6. Failing to capture these speaker characteristics may lead to missing context, misaligned entity associations, and, ultimately, incorrect factual summaries. The significance of examining speaker dynamics and roles is well-recognized in other fields such as linguistics (?) and is, together with the language challenge of the preceded analysis, a common hook to demonstrate and motivate techniques for dialogue summarization.
@@
Figure 6: Example conversation of the speaker challenge with defined roles (Alex as customer, Jordan as customer support agent, Taylor as technical specialist). The dialogue shows a change of roles when Jordan calls Taylor (marked with orange), shifting from Jordan as a primary troubleshooter to a facilitator for Taylor’s technical interventions. Further, multiple direct references to entities highlighted with blue (e.g., ”device”, ”error codes”) and indirect references to participants in red (e.g., ”you”) are incorporated.
@@
Approaches. The speaker challenge arises from identifying participants, tracking their actions and entity references, and revealing participant relationships. These challenges can be aligned with the difficulties of named entity recognition (NER), coreference resolution, and dependency parsing. Although advancements with Transformer-based models regarding NER, the explored techniques struggle to accurately identify and categorize entities across different text types and domains and with unstructured texts (?, ?, ?), rendering realistic dialogue scenarios as a challenging field. Coreference resolution poses difficulties for techniques in accurately identifying mentions and contextual links (?) and understanding the linguistic structures and relationships (?), extending the understanding challenge (Section 3.3).
@@
Works typically use dense vectors (?, ?) to accurately represent speaker roles in dialogue and act as labels for utterances (?, ?), offering insights into each speaker’s function within the dialogue. Speaker vectors are either randomly initialized and trainable (?, ?) or produced by small neural networks (?, ?) and represent different roles, e.g., ”industrial designer” (?), ”judge” (?), which are appended to the embedding of the speaker’s turn. For capturing speaker dynamics and resolving coreferences, graph-based models are an established approach, e.g., to build a graph representing each speaker’s main ideas and their discourse, alongside their interactions, reflecting inner- and inter-speaker structures (?, ?).
@@ 3.5 Salience
Characteristics. The salience challenge involves identifying and summarizing the most relevant information within a dialogue (?). This includes identifying content critical for insightful summaries (?), where the understanding of critical shifted lately from content for general audience summaries to summaries aligning with the viewpoints, values, and ethics of specific participant (?), and locating text spans containing relevant content (?). Figure 7 demonstrates the differences between personalized and generally applicable summaries, as the generated texts for Simon and Nadia are tailored to their role, stands, and tasks, whereas the general summary provides a neutral overview. Failure to effectively locate the scattered information or understand what is considered salient can impact the depth and relevance of the generated summaries, focusing on unimportant content while omitting relevant ones (?, ?, ?, ?).
@@
To nest the general understanding of what is considered salient into models, established strategies include additional training stages to better distinguish between salient and non-salient content. ? (?) introduce the uncovered loss to point out if salient information is missing and the contrastive loss to distinguish between relevant and irrelevant sentences, respectively. Negative samples, categorized as redundant (i.e., text with unnecessary utterances) and null (i.e., text with relevant utterances removed), help models prioritize important content (?, ?). To better determine salient content for specific participants, works propose a fine-tuning task in which the topic or user perspective to be focused on is passed as an input along the dialogue text (?), apply question-driven pre-training (?) and use dynamic prompts to direct model attention to key dialogue aspects (?). Additional strategies involve direct text manipulation (?), personal named entity planning to concentrate on specific entities (?), adapting attention mechanisms (?), and integrating human feedback (?) or graph structures (?) to further tailor summaries to the dialogue’s essential parts.
@@ 3.6 Factuality
In contrast to the previous input-related challenges, the factuality challenge describes the problem of correcting false content in a generated summary caused by a non-robust architecture. Factual incorrect content contains intrinsic errors, such as hallucinating events not present in the original transcript, and extrinsic errors, such as incorrect coreference resolution misrepresenting actual event details, and wrong conclusions stemming from negation (?). Intrinsic and extrinsic errors are shown in a sample dialogue in Figure 8. Given the importance of ensuring dialogue summarizers accurately reflect the conversation, research focuses on creating safety measures to maintain summary credibility despite these challenges.
Figure 8: Example of the factuality challenge where the generated summary exhibits intrinsic errors through hallucinating events (blue background) such as the finalization of a marketing strategy and a detailed review of financial strategies. The extrinsic error reporting the wrong date of the meeting is marked with red.
Approaches. Research interest predominantly targets post-processing approaches that rewrite the generated summary to correct factual errors. Notable approaches here are the usage of LLMs for error detection and correcting (?) and reformulating the summary to closely match the original dialogue’s predictive capabilities for subsequent content (?). Other approaches explore the use of auxiliary tasks (?) to estimate the factual aspects and predict the missing aspects in a summary (?), losses such as encouraging a model to generate sentences about content not yet covered and differentiating factual from non-factual sentences (?, ?), and human feedback (?). Further, architecture-modifying approaches are explored, such as a new encoder to grasp dialogue states (?) or to handle graph structures, where events of dialogue are captured and organized in a graph. A special attention mechanism is used to fill in summary (?), and a hierarchical approach builds on sub-summaries (?). Another noteworthy technique is the generation of a sketch to structure out the final summary along this plan (?). Negation is a long-standing challenge for language models often neglected in computational studies (?, ?).
@@ 4 Datasets
In this review, we collected both methods and downstream datasets from the selected publications. We only consider datasets used at least five times in our analysis to avoid including resources with little community impact. This yields 18 datasets that we group into six categories based on the dialogue subdomains they stem from: daily chat, online chat, meeting, screenplay, customer service, and medical. We also create another category for datasets that don’t fit neatly into these subdomain groups.
These datasets typically condense a short dialogue of about 1k tokens between two real persons into a third-person summary of about 100 tokens. More complex scenarios, such as meetings, parliamentary debates, and TV series dialogues, involve around 10k to 20k tokens in the transcript and five or more participants. Due to privacy concerns, a significant portion of currently publicly available datasets are reenactments based on real events, e.g., by Amazon Mechanical Turk workers, or established datasets such as SAMSum (?) and AMI (?) are modified to generate new data, e.g., by changing the dialogue transcript to a third-person report (?). This underscores that the variety of existing datasets is comparably limited, and these datasets may not capture real-life scenarios. The 18 available established datasets are detailed in Sections 4.1, 4.2, 4.3, 4.5, 4.7, 4.6, 4.4 and 4.8. Given the limited number of available datasets, we present methods for creating artificial datasets and strategies for optimizing the use of existing datasets in Section 4.9.
For this section, we consider 93 papers to identify the established datasets, excluding works that do not mention a dataset, datasets that are no longer publicly available, and datasets in a language other than English. We did not assess dataset quality or perform a detailed analysis of the inherent challenges, leaving that for future research. In Figure 9, we connect the datasets with their respective primary challenge, as noted by their creators, to provide an overview of the distribution of challenges in datasets.
Figure 9: Matching between established datasets and their primary reported challenge.

### v3

@@ 3.1 Language
Characteristics. The language challenge describes the idiosyncratic nature of spoken language and individual speech patterns. It includes linguistic subtleties such as informal expressions (e.g., ‘yeah,’), ungrammatical structures, colloquialisms, personal vocabulary, and linguistic noise such as filler words (?, ?, ?, ?, ?). It further covers content repetition, i.e., speakers restate or rephrase information for emphasis or clarity (?, ?, ?) and domain-specific terminology, e.g., medical terms (?, ?, ?, ?). These elements demand that models accurately interpret and adapt to the linguistic features of dialogues (?). Figure 3 presents a dialogue snippet showing examples of these three sub-challenges. Failing to handle this challenge may result in a loss of coherence and clarity, repetition of content, and factuality issues.
Figure 3: Dialogue snippet showing examples of the idiosyncratic nature of spoken language and individual speech patterns. Red displays disfluencies, blue highlights personal speech patterns, orange stands for colloquialism, green represent informal expressions.
Approaches. Transformer-based models often struggle with the language characteristics due to a gap between their pre-training data, typically well-edited texts (e.g., news articles, research papers, Wikipedia entries, ?, 2020), and the characteristics of spoken dialogue not reflected in these edited texts (?). The transfer between these text styles is hindered by the scarcity of diverse, dialogue-oriented pre-training data, resulting in a lack of exposure to dialogue data during training, consequently reducing models’ performance. Repetition of already provided information is less of an issue for current language models (?).
Current research aims to bridge this gap between the formal language of the pre-training data and dialogues by retraining models with dialogue-focused tasks, considering both single- and multi-task setups. Explored tasks span masking key dialogue elements (e.g., pronouns entities, high-content tokens, and words, ?, 2021) and part-of-speech tagging (?). More recent tasks aim to transform informal, first-person dialogue into a structured, third-person narrative through changing speaker names, adjusting grammar, and adding emotional context (?), or simulating a dialogue by modifying pre-training documents into a conversation structure (?). Teaching models to understand the ‘who-did-what’ structure through a specific task is further gaining traction (?). Other explored strategies include pre-processing the input for anaphora resolution (?) and adapted pre-training, e.g., to adjust to specific terminology (?, ?, ?).
@@ 3.2 Structure
Characteristics. Aiming to extract a conversation’s built-up, the structure challenge is about segmenting a dialogue transcript and creating an ordered summary based on the extracted structures (?). This challenge involves identifying the topic flow of a conversation (?, ?, ?, ?), tracking the dialogue phases (e.g., problem identification, decision making, ?, 2021; ?, 2023), and analyzing the utterance dependency (i.e., relationship and dependencies between utterances, ?, 2021a; ? 2021b; ?, 2022). The dialogue in Figure 4 shows topic flow, dialogue phases, and utterance dependency. This challenge can lead to summaries lacking coherence, completeness, and depth (?, ?, ?, ?, ?).
@@
Approaches. Transformer-based models face two primary hurdles when handling the structural challenge. First, they must adapt to the variability in dialogue structures, such as differing topic sequences and conversation phases depending on the conversation type. Second, they need to track and connect long-range dependencies to link utterances. The first difficulty is related to the lack of varied datasets across different dialogue types on which models can train to improve generalization as well as that most pre-trained models used are expecting more precise structural signals due to the more structured data used during pre-training (Section 3.1). The second problem relates to a weakness of transformed-based architectures: their limitation in handling long sequences, which stems from the practical limitation of a fix-sized context window (?), the blurred attention on a single token when many tokens are processed (?), and the bias towards recent information (?).
Recent methods employ graph structures to understand and use dialogue structures more effectively. They combine static and dynamic graphs for a detailed examination of conversation dynamics (?) where the static graphs represent unchanging aspects like speaker relationships, and dynamic graphs track how dialogues evolve. Abstract Meaning Representation (AMR) graphs are employed to capture overarching themes (?), detailed sentence-level connections, and entity interactions, enhancing content comprehension (?). Meanwhile, a hybrid approach combining traditional language models with graph neural networks captures the utterance dependency and addresses the long-range dependency sub-challenge (?). Similarly, ConceptNet (?) is used for thematic linking (?).
To manage long dialogue transcripts beyond context-size limits and the inherent long-distance relations, researchers are moving away from simply truncating inputs to context length to avoid losing vital information. An established approach segments the summarization process into stages (e.g., topics and dialogue phases) for coherent processing (?, ?, ?). The segmentation and multi-pass strategy iteratively summarizes parts of the dialogue and combines partial summaries with a subset of the remaining content in each iteration to build a complete summary (?). The $Summ^{N}$ framework (?) summarizes dialogue segments independently before refining them together, offering a scalable solution.
Other techniques involve additional training tasks (?, ?, ?, ?), sentence importance measures to group sentences into sub-topics (?, ?), while hierarchical setups to differentiate dialogue parts based on who is speaking (word level) and what is being discussed (turn level) are also popular (?).
Figure 5: Example conversation to demonstrate the comprehension challenge including direct and implied content. Directly stated (orange) are the explicit updates by Bob and Charlie about their respective tasks. Green marks implied content such as a deadline for a demo and a current focus on performance and finalizing features (unmentioned background knowledge), Redis implies an understanding of its role in performance enhancement (organizational knowledge and reference to prior discussions on data requirements).
@@ 3.3 Comprehension
Characteristics. The comprehension challenge involves accurately understanding and contextualizing utterances such that models can grasp and leverage these insights for summarization (?). This includes direct and implied content, e.g., unmentioned background knowledge, organizational knowledge, and prior discussions not explicitly referred to (examples in Figure 5). To infer and understand this content, the whole context of a dialogue is required, spanning both the local context, focusing on neighboring words and sentences, and the global context, covering the entire dialogue’s informational flow (?). Insufficient comprehension can lead to misunderstanding the content, omitting parts, or falsely presenting the information.
Approaches. Transformer-based models struggle with comprehension and contextualization due to the underlying dependency parsing challenges, particularly with unclear sentence boundaries as they appear in spoken language, which can lead to catastrophic forgetting and result in unreliability in handling long-range dependencies (?). Additionally, these models interpret the text literally, affecting their ability to grasp nuanced meanings (?, ?). As these models often rely on syntactic relationships between words, they miss the nuanced understanding humans use to interpret implied meanings of sarcasm, irony, euphemisms, and similar pragmatic elements (?). Consequently, this can lead to summaries that miss deeper meanings or implications, resulting in content that may be shallow or misleading (?).
@@ 3.4 Speaker
Characteristics. Participants, their interactions, and how they influence the dynamics of a conversation are part of the speaker challenge. This includes the individual participant roles (e.g., ‘project manager,’ ‘customer service agent’, ?, 2021b; ?, 2022), topic-dependent role changes (?, 2021; ?, 2022a; ?, 2022; ?, 2022; ?, 2022; ?, 2022; ?, 2023), their interactions (e.g., contributions depending on viewpoints, ?, 2021; ?, 2022; ?, 2022; and on backgrounds, ?, 2021; ?, 2021; ?, 2022), and references to entities and other participants (?, ?, ?, ?). Examples of the speaker characteristics are shown in Figure 6. Failing to capture these speaker characteristics may lead to missing context, misaligned entity associations, and, ultimately, incorrect factual summaries. The significance of examining speaker dynamics and roles is well-recognized in other fields such as linguistics (?) and is, together with the language challenge of the preceded analysis (Section 3.1), a typical hook to demonstrate and motivate techniques for dialogue summarization.
@@
Figure 6: Example conversation of the speaker challenge with defined roles (Alex as customer, Jordan as customer support agent, Taylor as technical specialist). The dialogue shows a change of roles when Jordan calls Taylor (marked with orange), shifting from Jordan as a primary troubleshooter to a facilitator for Taylor’s technical interventions. Further, multiple direct references to entities highlighted with blue (e.g., ‘device’, ‘error codes’) and indirect references to participants in red (e.g., ‘you’) are incorporated.
@@
Approaches. The speaker challenge stems from identifying participants, tracking their actions and entity references, and revealing participant relationships. These aspects correspond to the difficulties of named entity recognition (NER), coreference resolution, and dependency parsing. Despite advancements with Transformer-based models regarding NER, current techniques struggle with entity identification and categorization, particularly in unstructured texts and across diverse domains (?, ?, ?). This limitation is especially pronounced in realistic dialogue scenarios. Coreference resolution poses difficulties for techniques in accurately identifying mentions and contextual links (?) and understanding the linguistic structures (?), extending the understanding challenge (Section 3.3).
@@
Works typically use dense vectors (?, ?) to accurately represent speaker roles in dialogue and act as labels for utterances (?, ?), offering insights into each speaker’s function within the dialogue. Speaker vectors are either randomly initialized and trainable (?, ?) or produced by small neural networks (?, ?) and represent different roles, e.g., ‘industrial designer’ (?), ‘judge’ (?), which are appended to the embedding of the speaker’s turn. For capturing speaker dynamics and resolving coreferences, graph-based models are an established approach, e.g., representing each speaker’s main ideas and their discourse, alongside their interactions, reflecting inner- and inter-speaker structures (?, ?).
@@
Figure 8: Example of the factuality challenge where the generated summary exhibits extrinsic errors through hallucinating events (blue background) such as the finalization of a marketing strategy and a detailed review of financial strategies. The intrinsic error reporting the wrong date of the meeting is marked with red.
@@ 3.5 Salience
Characteristics. The salience challenge involves identifying and summarizing the most relevant information within a dialogue (?). This includes identifying content critical for insightful summaries (?), where the understanding of critical shifted lately from content for general audience summaries to summaries aligning with the viewpoints, values, and ethics of specific participant (?), and locating text spans containing relevant content (?). Figure 7 demonstrates the differences between personalized and generally applicable summaries, as the generated texts for Simon and Nadia are tailored to their role, stands, and tasks, whereas the general summary provides a neutral overview. In contrast to traditional summarization, where salience may be related to the position of sentences (positional salience), dialogues are more dynamic, and salience emerges from how ideas evolve over the conversation. Failure to effectively locate the scattered information or understand what is considered salient can impact the depth and relevance of the generated summaries, focusing on unimportant content while omitting relevant information bits (?, ?, ?, ?).
@@
To nest the general understanding of what is considered salient into models, established strategies include additional training stages to better distinguish between salient and non-salient content. ? (?) introduce the uncovered loss to point out if salient information is missing and the contrastive loss to distinguish between relevant and irrelevant sentences, respectively. Negative samples, categorized as redundant (i.e., text with unnecessary utterances) and null (i.e., text with relevant utterances removed), help models prioritize important content (?, ?). To better determine salient content for specific participants, works propose a fine-tuning task in which the topic or user perspective to be focused on is passed as an input along the dialogue text (?), apply question-driven pre-training (?) and use dynamic prompts to direct model attention to key dialogue aspects (?). Additional strategies involve direct text manipulation (?), personal named entity planning to concentrate on specific entities (?), adapting attention mechanisms (?), and integrating human feedback (?) or graph structures (?) to further tailor summaries to the dialogue’s core information.
Challenge Related Errors
Language incoherence, repetition, hallucination
Structure incoherence, omission, lack of depth
Comprehension lack of context understanding, omission, hallucination
Speaker lack of context understanding, false coreference resolution, hallucination
Salience irrelevance, lack of context understanding
Factuality hallucination
Table 2: Definition of eight to-be-expected error types in dialogue summarization, based on existing meeting summarization related error types (?).
@@ 3.6 Factuality
In contrast to the previous input-related challenges, the factuality challenge describes the problem of correcting false content in a generated summary caused by a non-robust architecture. Factual incorrect content contains extrinsic errors, such as hallucinating events not present in the original transcript, and intrinsic errors, such as incorrect coreference resolution misrepresenting actual event details, and wrong conclusions stemming from negation (?). Extrinsic and intrinsic errors are shown in a sample dialogue in Figure 8. Given the importance of ensuring summaries accurately reflect the conversation, research focuses on creating safety measures to maintain credibility.
Approaches. Research interest predominantly targets post-processing approaches that rewrite the generated summary to correct factual errors. Notable approaches are the usage of LLMs for error detection and correcting (?) and reformulating the summary to closely match the original dialogue’s predictive capabilities for subsequent content (?). Other approaches explore the use of auxiliary tasks (?) to estimate the factual aspects and predict the missing aspects in summary (?), losses such as encouraging a model to generate sentences about content not yet covered and differentiating factual from non-factual sentences (?, ?), and human feedback (?). Further, architecture-modifying approaches are explored, such as a new encoder to grasp dialogue states (?) or to handle graph structures, where dialogue events are captured and organized in a graph. A slot-driven beam search algorithm is used in a filling-in-the-blanks setup to give priority to generating salient elements in the summary (?), and a hierarchical approach builds on sub-summaries (?). Another noteworthy technique is the generation of a sketch to structure the final summary along this plan (?). Negation remains a long-standing challenge for language models often neglected in computational studies (?, ?).
@@ 4 Datasets
In this review, we collected data generation methods and downstream datasets from the selected publications. We only consider datasets used at least five times in our analysis to avoid including resources with little community impact. This setup yields 18 datasets that we group into six categories based on the dialogue subdomains they stem from: daily chat, online chat, meeting, screenplay, customer service, medical, and debates. We also create an ‘others’ category for datasets that do not fit neatly into these subdomain groups.
These datasets typically condense a short dialogue of about 1k tokens between natural persons into a third-person summary of about 100 tokens. More complex scenarios, such as meetings, parliamentary debates, and TV series dialogues, involve around 10k to 20k tokens in the transcript and at least four participants. Due to privacy concerns, a significant portion of currently publicly available datasets are reenactments based on actual events, e.g., by Amazon Mechanical Turk [^7] https://www.mturk.com/ workers, or established datasets such as SAMSum (?) are modified to generate new data, e.g., by changing the dialogue transcript to a third-person report (?). This underscores that the variety of existing datasets is comparably limited, and these datasets may not capture real-life scenarios. The 18 available established datasets are detailed in Sections 4.1, 4.2, 4.3, 4.5, 4.7, 4.6, 4.4 and 4.8. Given the limited number of available datasets, we present methods for creating artificial datasets and strategies for optimizing the use of existing datasets in Section 4.9.
For this section, we consider 93 papers to identify the established datasets, excluding works that do not mention a dataset, datasets that are no longer publicly available, and datasets in a language other than English. We did not assess dataset quality or perform a detailed analysis of the inherent challenges, leaving that for future research. In Table 3, we connect the datasets with their respective primary challenge, as noted by their creators, to provide an overview of the distribution of challenges in datasets.